An electric heater is rated at 1500 W, a toaster at 750 W, and an electric grill at 1000 W. The three appliances are connected in parallel to common 120 V household circuit. (a) How much current does each draw? (b) Is a circuit with a 25.0 Amp circuit breaker sufficient in this situation? Numerically justify your answer.
Answer #1
.
Problem statement: An electric heater is rated at 1500W, a toaster at 750 W, and an electric grill at 1000 W. Thethree appliances are connected in parallel to common 120 Vhousehold circuit. (a) How much current does eachdraw? (b) Is a circuit with a 25.0 Amp circuit breakersufficient in this situation? Numerically justify youranswer.
.
01.   Power =(voltage)*(current)      ∴      Current= (Power)/(voltage)
02.   Iheater = (1500 Watts)/(120 volts)= 12.50 Ampere
03.   Itoaster = (750 Watts)/(120 volts)= 6.25 Ampere
04.   Igrill = (1000 Watts)/(120 volts) =8.33 Ampere
05.   Itotal = Iheater +Itoaster + Igrill = 12.5 A + 6.25 A + 8.33 A= 27.08 Ampere
06.   No, the load exceeds the circuitbreaker's value by more than 2 Ampere.
.